OpenStudy (anonymous):

The current in a wire is 0.72 A. Calculate the charge that passes through the wire in 7 minutes

OpenStudy (anonymous):

@aayushi.somani

OpenStudy (aakashsudhakar):

The definition of current is given as the average change of charge over time, also given as... \[I = \frac{ dQ }{ dt } \approx \frac{ \Delta Q }{ \Delta t }\]Using this, we can approximate the total charge that passed through the wire to be equal to the product of the total current and the total time elapsed, also given by... \[\Delta Q \approx I \times \Delta t\]This gives you an equation that you can simply plug in the given values of the total current (I = 0.72A) and the total elapsed time (t = 420sec). Solving gets you that Q = (I)(t) = (0.72A)(420sec) = 302.4 Coulombs.

OpenStudy (anonymous):

thanks